# Enhancing Boundary Segmentation for Topological Accuracy with Skeleton-based Methods

### v1

@@ Abstract
Topological consistency plays a crucial role in the task of boundary segmentation for reticular images, such as cell membrane segmentation in neuron electron microscopic images, grain boundary segmentation in material microscopic images and road segmentation in aerial images. In these fields, topological changes in segmentation results have a serious impact on the downstream tasks, which can even exceed the misalignment of the boundary itself. To enhance the topology accuracy in segmentation results, we propose the Skea-Topo Aware loss, which is a novel loss function that takes into account the shape of each object and topological significance of the pixels. It consists of two components. First, the skeleton-aware weighted loss improves the segmentation accuracy by better modeling the object geometry with skeletons. Second, a boundary rectified term effectively identifies and emphasizes topological critical pixels in the prediction errors using both foreground and background skeletons in the ground truth and predictions. Experiments prove that our method improves topological consistency by up to 7 points in VI compared to 13 state-of-art methods, based on objective and subjective assessments across three different boundary segmentation datasets. The code is available at https://github.com/clovermini/Skea_topo.
@@ 3.1 Skeleton Aware Weighted loss
The design of our loss is based on two principles. First, for the boundary, higher weights should be assigned to pixels in narrower boundaries, since they have a higher risk of fracture. Second, for the object, higher weights are assigned to misclassified pixels far away from the boundary, since those near it can be difficult to distinguish, often due to boundary ambiguity and manual annotation misalignment. Existing weighting methods use the distance transform function to assign weights to individual objects. For the stability of training, these methods normalize the distance value using the maximum distance, so relatively low weights are assigned to narrow areas of irregular objects, which is prone to boundary closure errors. To overcome this limitation, our method introduces the object skeleton for distance calculation and weight normalization. The skeleton of an object is extracted through the morphological skeletonization function provided by scikit-imagevan der Walt et al. (2014). Through the object skeleton, a uniform weight distribution can be achieved in narrow areas and other areas, as shown in Fig. 2 (c). The detailed definitions of $w_{1}^{s}(x)$ and $w_{0}^{s}(x)$ are as follows:
@@ 3.2 Boundary Rectified Term
Skeletons are used because they can effectively exclude most of the $ff$ near the boundary but not areas that significantly alter the shape or are located far from the boundary. Such errors should also be assigned a heavier penalty to maintain the accuracy of the results. Instead of individually examining each pixel to determine whether flipping that pixel will change the topology as WarpHu (2022), our approach identifies all critical pixels using a few image-level operation steps through the skeletons. The most time-intensive aspect of this process is computing the foreground and background skeletons, each with a time complexity of O(n), where n represents the number of image pixels. Leveraging parallelization and hardware acceleration technologies in Python, our method ensures fast execution speeds in practical applications.
@@ 4.1.3 Evaluation Metrics
Five metrics were used in the experiment to evaluate the performance of each method. These metrics include a pixel-wise metric (Dice) and topology-related metrics: Variation of Information (VI) Nunez-Iglesias et al. (2013), Adjusted Rand Index (ARI) Vinh et al. (2009), mean average precision (mAP) Lin et al. (2014), and Betti Error (Betti). For detailed descriptions of these metrics, please refer to the Appendix. VI is given prominence in our evaluation because it is sensitive to the presence of topological changes and the extent of their impact. To ensure fair evaluation, post-processing was applied to the segmentation results and ground truth. It includes three steps: boundary dilation with a square filter of size 3 to remove small holes, skeletonization, and a second round of dilation using the same filter.
@@ 4.2 Qualitative and Quantitative Results
Table 1 displays the quantitative results of our loss function compared to 14 other SOTA losses. Our method delivers top-tier performance across most metrics: it ranks first in most cases and a close second in some cases. Other methods exhibit their own strengths and weaknesses on different metrics, whereas our approach excels across all evaluated metrics, which showcases its well-rounded effectiveness. Compared to the existing methods, our approach achieves a maximum improvement of 7 points in VI. Significantly, no other method has simultaneously shown superior performance across all three datasets as ours does, which underscores the exceptional effectiveness and robustness of our approach. The subscript data with $\pm$ represent the std of the results of three-fold experiments. The relatively low std of our method highlights its stability. All reported metrics of our method are based on the best overall performance in the ablation studies. Figs. 3, 4, and 5 qualitatively compare the boundary segmentation results between our method and other methods across three datasets. These figures demonstrate that our method is more effective than other methods in eliminating topological errors when compared to the ground truth (GT).
@@ 4.3 Ablation Studies
In this section, only the ablation experiment results on the SNEMI3D dataset are reported, and the loss weight $\lambda$ was set to 1 by default, which already significantly improves performance. For more comprehensive results of the ablation experiments, please refer to the Appendix.
@@ 4.3.2 Impact of the loss terms
Ablation experiments on each component of our proposed loss were performed to verify their necessity. Specifically, the components of the penalty term in BoRT were evaluated, w/o $ff$ refers to removing the $ffp$ and $ffn$ terms in $L_{bort\_topo}$, and w/o $ff$ indicates removing the $tn$ and $tp$ terms. The results in Table 2 demonstrate the necessity of all penalty components. Furthermore, similar ablation studies were conducted on the other two datasets. The IRON and SNEMI3D datasets have generally consistent conclusions, but the standard form of the penalty term poorly performs on the MASS. ROAD dataset, possibly due to the extreme data imbalance in the MASS. ROAD dataset. Consequently, we modified the approach by weighting $tfn$ and removing the $ff$ component. For a detailed discussion and results of these experiments, please refer to the Appendix.
@@ 5 Conclusion
In this work, a Skea-Topo Aware loss was developed to guide the boundary segmentation network towards generating topologically consistent results. It includes a skeleton aware weighted loss, which better models geometric information, and a boundary rectification term, which penalizes topological critical pixels. To efficiently identify the critical pixels, a method based on the foreground and background skeletons was proposed. The proposed method was thoroughly evaluated through rigorous comparative and ablation experiments on three datasets, and it demonstrates a maximum improvement of 7 points in the VI index.

### v2

@@ Abstract
Topological consistency plays a crucial role in the task of boundary segmentation for reticular images, such as cell membrane segmentation in neuron electron microscopic images, grain boundary segmentation in material microscopic images and road segmentation in aerial images. In these fields, topological changes in segmentation results have a serious impact on the downstream tasks, which can even exceed the misalignment of the boundary itself. To enhance the topology accuracy in segmentation results, we propose the Skea-Topo Aware loss, which is a novel loss function that takes into account the shape of each object and topological significance of the pixels. It consists of two components. First, a skeleton-aware weighted loss improves the segmentation accuracy by better modeling the object geometry with skeletons. Second, a boundary rectified term effectively identifies and emphasizes topological critical pixels in the prediction errors using both foreground and background skeletons in the ground truth and predictions. Experiments prove that our method improves topological consistency by up to 7 points in VI compared to 13 state-of-art methods, based on objective and subjective assessments across three different boundary segmentation datasets. The code is available at https://github.com/clovermini/Skea_topo.
@@ 3.1 Skeleton Aware Weighted loss
The design of our loss is based on two principles. First, for the boundary, higher weights should be assigned to pixels in narrower boundaries, since they have a higher risk of fracture. Second, for the object, higher weights are assigned to misclassified pixels far away from the boundary, since those near it can be difficult to distinguish, often due to boundary ambiguity and manual annotation misalignment. Existing weighting methods use the distance transform function to assign weights to individual objects. For the stability of training, these methods normalize the distance value using the maximum distance, so relatively low weights are assigned to narrow areas of irregular objects, which are prone to boundary closure errors. To overcome this limitation, our method introduces the object skeleton for distance calculation and weight normalization. The skeleton of an object is extracted through the morphological skeletonization function provided by scikit-imagevan der Walt et al. (2014). Through the object skeleton, a uniform weight distribution can be achieved in narrow areas and other areas, as shown in Fig. 2 (c). The detailed definitions of $w_{1}^{s}(x)$ and $w_{0}^{s}(x)$ are as follows:
@@ 3.2 Boundary Rectified Term
Skeletons are used because they can effectively exclude most of the $ff$ near the boundary but not areas that significantly alter the shape or are located far from the boundary. Such errors should also be assigned a heavier penalty to maintain the accuracy of the results. Instead of individually examining each pixel to determine whether flipping that pixel will change the topology as WarpHu (2022), our approach identifies all critical pixels using a few image-level operation steps through the skeletons. The most time-intensive step of this process is computing the foreground and background skeletons, each with a time complexity of O(n), where n represents the number of image pixels. Leveraging parallelization and hardware acceleration technologies in Python, our method ensures fast execution speeds in practical applications.
@@ 4.1.3 Evaluation Metrics
Five metrics were used in the experiment to evaluate the performance of each method. These metrics include a pixel-wise metric (Dice) and topology-related metrics: Variation of Information (VI) Nunez-Iglesias et al. (2013), Adjusted Rand Index (ARI) Vinh et al. (2009), mean average precision (mAP) Lin et al. (2014), and Betti Error (Betti). VI is given prominence in our evaluation because it is sensitive to the presence of topological changes and the extent of their impact. To ensure fair evaluation, post-processing was applied to the segmentation results and ground truth. It includes three steps: boundary dilation with a square filter of size 3 to remove small holes, skeletonization, and a second round of dilation using the same filter.
@@ 4.2 Qualitative and Quantitative Results
Table 1 displays the quantitative results of our loss function compared to 13 other SOTA losses. Our method delivers top-tier performance across most metrics: it ranks first in most cases and a close second in some cases. Other methods exhibit their own strengths and weaknesses on different metrics, whereas our approach excels across all evaluated metrics, which showcases its well-rounded effectiveness. Compared to the existing methods, our approach achieves a maximum improvement of 7 points in VI. Significantly, no other method has simultaneously shown superior performance across all three datasets as ours does, which underscores the exceptional effectiveness and robustness of our approach. The subscript data with $\pm$ represent the std of the results of three-fold experiments. The relatively low std of our method highlights its stability. All reported metrics of our method are based on the best overall performance in the ablation studies. Figs. 3, 4, and 5 qualitatively compare the boundary segmentation results between our method and other methods across three datasets. These figures demonstrate that our method is more effective than other methods in eliminating topological errors when compared to the ground truth (GT).
@@ 4.3 Ablation Studies
In this section, only the ablation experiment results on the SNEMI3D dataset are reported, and the loss weight $\lambda$ was set to 1 by default, which already significantly improves performance.
@@ 4.3.2 Impact of the loss terms
Ablation experiments on each component of our proposed loss were performed to verify their necessity. Specifically, the components of the penalty term in BoRT were evaluated, w/o $ff$ refers to removing the $ffp$ and $ffn$ terms in $L_{bort\_topo}$, and w/o $ff$ indicates removing the $tn$ and $tp$ terms. The results in Table 2 demonstrate the necessity of all penalty components. Furthermore, similar ablation studies were conducted on the other two datasets. The IRON and SNEMI3D datasets have generally consistent conclusions, but the standard form of the penalty term poorly performs on the MASS. ROAD dataset, possibly due to the extreme data imbalance in the MASS. ROAD dataset. Consequently, we modified the approach by weighting $tfn$ and removing the $ff$ component.
@@ 5 Conclusion
In this study, a Skea-Topo Aware loss was developed, tailored specifically to guide boundary segmentation networks towards the generation of topologically consistent results. It comprises a skeleton-aware weighted loss, which is enhanced to model geometric information more precisely by utilizing object skeletons, and a boundary rectification term, which imposes penalties on topologically critical pixels. These two components work together synergistically to improve segmentation performance. To efficiently identify the topologically critical pixels, a method based on the foreground and background skeletons was proposed. Our proposed method underwent rigorous evaluation through comparative and ablation experiments conducted on three datasets, demonstrating a maximum improvement of 7 points in the VI index, thereby validating its effectiveness and potential.